OpenStudy (anonymous):

In a group assignment, students are required to fill 10 beakers with 0.720 M CaCl2. If the molar mass of CaCl2 is 110.98 g/mol and each beaker must have 250. mL of solution, what mass of CaCl2 would be used? 3.13 g 38.5 g 200 g 617 g

OpenStudy (jfraser):

this is really an exercise in converting units. You know the volume that has to go in each beaker, you know the number of beakers that have to be filled, the concentration of the solution, and the molar mass of the calcium chloride. Each of those equalities can be turned into a conversion fraction so the units cancel and you're left with the simple math to find the answer\[10 \space beakers * \frac{0.250L}{1 \space beaker}*\frac{0.720mol\space CaCl_2 }{1L \space solution}* \frac{110.98g \space CaCl_2}{1\space mole \space CaCl_2}\]

OpenStudy (jfraser):

see how each unit is canceled and replaced by something else? until you get to the last unit, which is mass, which is what the question is asking